# In-context Time Series Predictor

### v1

@@ Abstract
Recent Transformer-based large language models (LLMs) demonstrate in-context learning ability to perform various functions based solely on the provided context, without updating model parameters. To fully utilize the in-context capabilities in time series forecasting (TSF) problems, unlike previous Transformer-based or LLM-based time series forecasting methods, we reformulate "time series forecasting tasks" as input tokens by constructing a series of (lookback, future) pairs within the tokens. This method aligns more closely with the inherent in-context mechanisms, and is more parameter-efficient without the need of using pre-trained LLM parameters. Furthermore, it addresses issues such as overfitting in existing Transformer-based TSF models, consistently achieving better performance across full-data, few-shot, and zero-shot settings compared to previous architectures.
@@ 1 Introduction
Transformer-based large language models (LLMs) have significantly impacted various research and application areas [1]. Their inherent in-context learning (ICL) capabilities highlighted by previous studies [2, 3, 4] allow them to adapt and generalize from context examples provided in input prompts without any parameter updates. This enables LLMs to effectively handle few-shot and zero-shot tasks. Recent research on ICL [5, 6, 7, 8, 9, 10] has shown that Transformers can adaptively learn to perform various functions including linear predictors, shallow MLPs, gradient descent, algorithm selection, etc., based on a series of (input, label) pairs as input tokens, as shown in Figure 1 (a)
@@
Time series forecasting (TSF) is critical in fields like epidemiology, finance, and transportation [11, 12, 13, 14], predicting future values from historical data. Temporal-wise Transformers, which build input tokens from series values at each timestep, have been widely researched in TSF [15, 16, 17, 18]. Some studies have identified several issues with such Transformers, like timestep mixing and permutation invariance [19], leading to overfitting and underperformance on real-world datasets compared to simpler models like linear predictors. Previously proposed solutions include channel independence [18], random channel dropout [20], and using Series-wise Transformers [21, 22], which consider each series as a token. Yet, their underlying mechanisms are not well explained. Additionally, the fixed structure of input series of these existing Transformers restricts their adaptability in few-shot and zero-shot learning scenarios for multivariate TSF.
Moreover, recent research have expanded the application of Transformer-based LLMs to TSF [23], achieving improvements in few-shot and zero-shot generalization. They use the methods like prompt engineering [24], fine-tuning and embedding inversion [25, 26] to integrate time series context into prompts, improving forecasting accuracy for new data. However, these methods are designed to adapt LLMs to TSF tasks rather than directly addressing the core aspects of TSF problems. This adaptation leads to inefficient use of LLM parameters and substantially increased computational costs.
@@
c) We show that the ICTSP structure encompasses several simpler models as special cases, allowing for a sequential adaptive reduction in complexity: i) predictors learned from context examples through ICL, ii) a series-wise Transformer without context examples [21], and iii) univariate MLPs or linear predictors [19]. This connection ensures stable performance across different complexities of time series dataset and prevents the significant overfitting that has previously hindered Transformers from outperforming simpler models consistently in real-world datasets.
@@ 2.1 Preliminaries
where $\mathbf{p}_{\cdot}$ are positional embeddings and $\mathbf{o}$ is a placeholder embedding for the target, typically zero-filled. Note that in some ICL settings, $\mathbf{x}_{\cdot}$ and $\mathbf{y}_{\cdot}$ may be divided into separate tokens, which is proved interchangeable with our expression above [10, 27]. The positional embeddings $\mathbf{p}_{\cdot}$ are added to the tokens and shown concatenated here just for illustration. Finally, the predicted output for the $i$-th data point, $\widehat{\mathbf{y}}_{i}$, can be extracted from the corresponding token in the transformer’s output $\mathbf{\hat{Z}}=\mathtt{TF}(\mathbf{Z})$. For simplicity, we have omitted the input and output linear projections for matrices $\mathbf{Z}$ and $\mathbf{\hat{Z}}$ in this description. In practice, these projections map the tokens to a $d$-dimensional latent space. The computation is formulated as $\mathbf{\hat{Z}}=\mathbf{W}_{\text{out}}\mathtt{TF}(\mathbf{W}_{\text{in}}\mathbf{Z}+\mathbf{b}_{\text{in}})+\mathbf{b}_{\text{out}}$, where $\mathbf{W}_{\text{in}}\in\mathbb{R}^{d\times(a+b)}$ and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{(a+b)\times d}$ are the linear layer weights, and $\mathbf{b}_{\text{in}}$ and $\mathbf{b}_{\text{out}}$ are the biases. The ICL settings allow the model to use $\{\mathbf{x}_{1},\mathbf{y}_{1},\cdots,\mathbf{x}_{i-1},\mathbf{y}_{i-1},\mathbf{x}_{i}\}$ to determine the best predictor $f_{i}^{*}:\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$.
@@ 2.2 Formulation of TSF Transformers
where $\mathbf{O}$ and $\{o_{1},\cdots,o_{C}\}$ represent zero-padding over the forecasting horizon. Again, the added positional embeddings $\mathbf{p}_{\cdot}$ are shown concatenated here just for illustration. Each token represents the time series channel structure at a specific timestep. Therefore, from the ICL perspective, the Transformer doesn’t learn the necessary prediction from $X_{I}$ to $X_{P}$ for TSF. Instead, it learns the mapping $f^{*}_{t}:\mathbf{p}_{t}\rightarrow[X^{(t)}_{1},\cdots,X^{(t)}_{C}]^{\top}$, which tries to find the underlying dynamic of the channel structure determined by the position from the context and applies it to the outputs. If true dependencies between the input series are lacking, which frequently occurs in real-world data, its focus on channel structure over temporal dependencies leads to significant overfitting. In §2.3, we will explain how existing methods like channel independence can partially address overfitting while introducing new issues. In the last part of §3, we will use examples to show that Temporal-wise Transformers perform well with strong channel dependencies but overfit when these dependencies are weak. Previous models using this architecture include LogTrans[15], Informer [16], Autoformer [17], etc. Note that while some of these models use an encoder-decoder structure, their token construction based on multi-channel values along the temporal dimension shares the same issues we highlight.
@@
Some recent methods [21, 22] transpose the input format above so each token represents an individual series, forming the Series-wise Transformer. This format, shown in Figure 2 (b), can be expressed as follows:
However, iTransformer [21] introduced this formulation believing it enhances modeling of inter-series dependencies compared to the Temporal-wise Transformer, which contrasts with our ICL analysis. Each token in the Series-wise Transformer represents an entire input series, capturing relationships between timesteps within that series. For series $j$, the Transformer learns a predictor $f^{*}_{j}:X_{j}^{(1:L_{I})}\rightarrow X_{j}^{(L_{I}+1:L_{I}+L_{P})}$, which also adjusts its own parameters to the inputs $\{X^{(1:L_{I})}_{i}\}_{i\neq j}$ from other series. Essentially, it acts more like a univariate predictor capable of adapting its parameters based on the context series. We show in the last part of §3 that the Series-wise Transformer is better suited for real-world datasets with weak multivariate relationships than those with strong inter-series dependencies, supporting our view from ICL. Also from this perspective, the Series-wise Transformer lacks ground truth context examples for the predictor, meaning it doesn’t fully utilize the Transformer’s ICL abilities. This limits its generalization and few-shot learning capabilities.
@@
Adaptive Model Reduction of ICTSP  Existing complex Temporal-wise Transformers often struggle to outperform simple baselines, like linear predictors, in noisy datasets lacking clear temporal-wise and channel-wise dependencies. In contrast, ICTSP adapts to different dataset complexities and reduces to simpler models when necessary. Specifically: i) When temporal-wise dependencies are weak, ICTSP can adaptively select simpler functions for TSF, such as linear predictors and shallow MLPs, as achievable by ICL. ii) When series patterns change greatly over time where the impact of context examples is weak, ICTSP can reduce to a Series-wise Transformer focusing only on local lookback of target tokens without context tokens. iii) When local lookback from other series have minimal contribution, ICTSP degrades to a simple univariate MLP, in which case the inter-token information in attention term from Eq.(1) disappears, and token $j$ at layer $k$ becomes $\mathbf{z}_{j,k}=\mathbf{z}_{j,k-1}+\mathtt{LN}(\mathtt{FFN}_{j,k}(\mathbf{z}_{j,k-1}))$. If $\mathtt{FFN}$ also becomes ineffective, ICTSP processes target tokens only with input and output linear projections outside the Transformer. At this stage, the model reduces to a linear predictor, effective for highly noisy time series data. ICTSP’s design provides appropriate shortcuts to enhance adaptability. We further explain in §A.4.1 how inappropriate shortcuts in the model reduction of the Temporal-wise Transformer limit its ability to learn dynamic patterns.
@@ 2.3 Solving Key Issues of TSF Transformers
PatchTST [18] suggested that univariate Temporal-wise Transformers, which inputs each channel independently (Figure 3 (b)), can address overfitting. This approach breaks the series-wise relationship in the tokens, transforming the ICL predictor into $f^{*}_{j,t}:\mathbf{p}_{t}\rightarrow X^{(t)}_{j}$, which predicts values solely from positions. This shift enhances the model’s focus on temporal dependencies while ignoring series-wise dependencies, making the formulation more suitable for real-world TSF tasks. Notably, the prompt structure used by recent LLMs for TSF, "… timestep: value …", also benefits from this approach.
ARM [20] introduced random dropping (Figure 3 (c)), which mitigates overfitting by randomly removing series from both input and output during training. This randomly breaks some relationships between channels within the tokens. With only positional information consistently preserved, the model focus more on the temporal generation process and retains only the most likely channel dependencies, reducing overfitting to non-existent channel relationships. However, the degree of channel dropout requires further tuning based on different dataset characteristics.
Methods such as LogTrans [15] and ARM [20] use strategies similar to Figure 3 (d) to build local representations for tokens across time intervals instead of single timesteps. This introduces temporal relationships within the token. For TSF data with weak channel dependencies, tokens naturally focus on intra-series temporal relationships (solid lines in Figure 3 (d)) rather than the weak channel-wise relationships (dashed lines in Figure 3 (d)) as in the original token format. This approach alleviates overfitting but requires careful tuning of the local window selection.
@@
Permutation Invariance   Studies like DLinear [19] and iTransformer [21] suggest that the poor performance of Temporal-wise Transformers is partly due to the inherent permutation invariance issue in Transformer, meaning swapping the positions of input tokens does not significantly affect the output. The Temporal-wise Transformer treats context tokens more as representations of the underlying channel structure. Thus, when a dataset lacks a temporally varying underlying channel structure, $\mathbf{p}_{t}$ in $f^{*}_{t}:\mathbf{p}_{t}\rightarrow[X^{(t)}_{1},\cdots,X^{(t)}_{C}]^{\top}$ becomes ineffective, and the output tends to converge to the context’s average. This is consistent with DLinear observations [19] where shuffling of input token minimally impact the output. Conversely, in the ICTSP, each context token represents a forecasting training example, so swapping their positions does not harm the forecasting. This makes the ICTSP better suited to the Transformer’s characteristics for TSF compared to previous structures.
Channel Structure Restriction  In the Temporal-wise Transformer, the number of input series and their relative positions within the tokens must be fixed. This limits the model’s ability to adapt to datasets with varying series numbers. Channel independence [18] may simplify the Temporal-wise Transformer to a univariate model, enhancing generalization but losing the ability to model inter-series dependencies. The Series-wise Transformer does not restrict the number of input series structurally but learns the characteristics of $C$ series during training through $\mathbf{p}_{j}$, making it hard to transfer embeddings to other datasets. Conversely, the ICTSP does not restrict the number of series and includes in-context training examples within the input, enabling easy transfer to input data of any size. Here, positional embeddings can be trained to only distinguish between samples without describing the characteristics of the series. The flexibility in input channel structure provides zero-shot multivariate forecasting ability, which previous multivariate models could not achieve.

### v2

@@ Abstract
Recent Transformer-based large language models (LLMs) demonstrate in-context learning ability to perform various functions based solely on the provided context, without updating model parameters. To fully utilize the in-context capabilities in time series forecasting (TSF) problems, unlike previous Transformer-based or LLM-based time series forecasting methods, we reformulate "time series forecasting tasks" as input tokens by constructing a series of (lookback, future) pairs within the tokens. This method aligns more closely with the inherent in-context mechanisms, and is more parameter-efficient without the need of using pre-trained LLM parameters. Furthermore, it addresses issues such as overfitting in existing Transformer-based TSF models, consistently achieving better performance across full-data, few-shot, and zero-shot settings compared to previous architectures  [^1] Code implementation is available at: https://github.com/LJC-FVNR/In-context-Time-Series-Predictor.
@@ 1 Introduction
Transformer-based large language models (LLMs) have significantly impacted various research and application areas (Brown et al., 2020). Their inherent in-context learning (ICL) capabilities highlighted by previous studies (Müller et al., 2022; Min et al., 2022; Wei et al., 2023; Xie et al., 2022) allow them to adapt and generalize from context examples provided in input prompts without any parameter updates. This enables LLMs to effectively handle few-shot and zero-shot tasks. Recent research on ICL (Zhang et al., 2023; Garg et al., 2022; Akyürek et al., 2023; Li et al., 2023; Dai et al., 2023; Bai et al., 2023) has shown that Transformers can adaptively learn to perform various functions including linear predictors, shallow MLPs, gradient descent, algorithm selection, etc., based on a series of (input, label) pairs as input tokens, as shown in Figure 1 (a).
@@
Time series forecasting (TSF) is critical in fields like epidemiology, finance, and traffic (Kaastra and Boyd, 1996; Lana et al., 2018; Yang et al., 2015; Ma et al., 2022), predicting future values from historical data. Temporal-wise Transformers, which build input tokens from series values at each timestep, have been widely researched in TSF (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Nie et al., 2022). Some studies have identified several issues with such Transformers, like timestep mixing and permutation invariance (Zeng et al., 2023), leading to overfitting and underperformance on real-world datasets compared to simpler models like linear predictors. Previously proposed solutions include channel independence (Nie et al., 2022), random channel dropout (Lu et al., 2024), and using Series-wise Transformers (Liu et al., 2024a; Wang et al., 2024), which consider each series as a token. Yet, their underlying mechanisms are not well explained. Additionally, the fixed structure of input series of these existing Transformers restricts their adaptability in few-shot and zero-shot learning for multivariate TSF.
Moreover, recent research have expanded the application of Transformer-based LLMs to TSF (Zhang et al., 2024), achieving improvements in few-shot and zero-shot generalization. They use the methods like prompt engineering (Gruver et al., 2023), fine-tuning and embedding inversion (Zhou et al., 2023; Jin et al., 2024) to integrate time series context into prompts, improving forecasting accuracy for new data. However, these methods are designed to adapt LLMs to TSF tasks rather than directly addressing the core aspects of TSF problems. This adaptation leads to inefficient use of LLM parameters and substantially increased computational costs.
@@
c) We show that the ICTSP structure encompasses several simpler models as special cases, allowing for a sequential adaptive reduction in complexity: i) predictors learned from context examples through ICL, ii) a series-wise Transformer without context examples (Liu et al., 2024a), and iii) univariate MLPs or linear predictors (Zeng et al., 2023). This connection ensures stable performance across different complexities of time series dataset and prevents the significant overfitting that has previously hindered Transformers from outperforming simpler models consistently in real-world datasets.
@@ 2.1 Preliminaries
$\mathbf{Z}_{k}=\mathtt{TF}_{k}(\mathbf{Z}_{k-1})=\mathbf{Z}_{k-1}+\mathtt{Attn}_{k}\left(\mathtt{LN}(\mathbf{Z}_{k-1})\right)+\mathtt{LN}\left(\mathtt{FFN}_{k}\left(\mathbf{Z}_{k-1}+\mathtt{Attn}_{k}\left(\mathtt{LN}(\mathbf{Z}_{k-1})\right)\right)\right),$
@@
$\mathbf{Z}^{(a+b)\times i}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{i}]=\begin{bmatrix}\mathbf{p}_{1}&\mathbf{p}_{2}&\ldots&\mathbf{p}_{i-1}&\mathbf{p}_{i}\\ \mathbf{x}_{1}&\mathbf{x}_{2}&\ldots&\mathbf{x}_{i-1}&\mathbf{x}_{i}\\ \mathbf{y}_{1}&\mathbf{y}_{2}&\ldots&\mathbf{y}_{i-1}&\mathbf{o}\end{bmatrix}$,
where $\mathbf{p}_{\cdot}$ are positional embeddings and $\mathbf{o}$ is a placeholder embedding for the target, typically zero-filled. Note that in some ICL settings, $\mathbf{x}_{\cdot}$ and $\mathbf{y}_{\cdot}$ may be divided into separate tokens, which is proved interchangeable with our expression above (Bai et al., 2023; Guo et al., 2024). The positional embeddings $\mathbf{p}_{\cdot}$ are added to the tokens and shown concatenated here just for illustration. Finally, the predicted output for the $i$-th data point, $\widehat{\mathbf{y}}_{i}$, can be extracted from the corresponding token in the transformer’s output $\mathbf{\hat{Z}}=\mathtt{TF}(\mathbf{Z})$. For simplicity, we have omitted the input and output linear projections for matrices $\mathbf{Z}$ and $\mathbf{\hat{Z}}$ in this description. In practice, these projections map the tokens to a $d$-dimensional latent space. The computation is formulated as $\mathbf{\hat{Z}}=\mathbf{W}_{\text{out}}\mathtt{TF}(\mathbf{W}_{\text{in}}\mathbf{Z}+\mathbf{b}_{\text{in}})+\mathbf{b}_{\text{out}}$, where $\mathbf{W}_{\text{in}}\in\mathbb{R}^{d\times(a+b)}$ and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{(a+b)\times d}$ are the linear layer weights, and $\mathbf{b}_{\text{in}}$ and $\mathbf{b}_{\text{out}}$ are the biases. The ICL settings allow the model to use $\{\mathbf{x}_{1},\mathbf{y}_{1},\cdots,\mathbf{x}_{i-1},\mathbf{y}_{i-1},\mathbf{x}_{i}\}$ to determine the best predictor $f_{i}^{*}:\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$.
@@ 2.2 Formulation of TSF Transformers
$\mathbf{Z}^{C\times L}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{L}]=\begin{bmatrix}\mathbf{P}_{1:L_{I}}&\mathbf{P}_{L_{I}+1:L_{I}+L_{P}}\\ X_{I}&\mathbf{O}\\ \end{bmatrix}=\begin{bmatrix}\mathbf{p}_{1}&\ldots&\mathbf{p}_{L_{I}}&\mathbf{p}_{L_{I}+1}&\ldots&\mathbf{p}_{L_{I}+L_{P}}\\ X_{1}^{(1)}&\ldots&X_{1}^{(L_{I})}&o_{1}&\ldots&o_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ X_{C}^{(1)}&\ldots&X_{C}^{(L_{I})}&o_{C}&\ldots&o_{C}\\ \end{bmatrix}$,
where $\mathbf{O}$ and $\{o_{1},\cdots,o_{C}\}$ represent zero-padding over the forecasting horizon. Again, the added positional embeddings $\mathbf{p}_{\cdot}$ are shown concatenated here just for illustration. Each token represents the time series channel structure at a specific timestep. Therefore, from the ICL perspective, the Transformer doesn’t learn the necessary prediction from $X_{I}$ to $X_{P}$ for TSF. Instead, it learns the mapping $f^{*}_{t}:\mathbf{p}_{t}\rightarrow[X^{(t)}_{1},\cdots,X^{(t)}_{C}$$^{\top}$, which tries to find the underlying dynamic of the channel structure determined by the position from the context and applies it to the outputs. If true dependencies between the input series are lacking, which frequently occurs in real-world data, its focus on channel structure over temporal dependencies leads to significant overfitting. In §2.3, we will explain how existing methods like channel independence can partially address overfitting while introducing new issues. In the last part of §3, we will use examples to show that Temporal-wise Transformers perform well with strong channel dependencies but overfit when these dependencies are weak. Previous models using this architecture include LogTrans(Li et al., 2019), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), etc. Note that while some of these models use an encoder-decoder structure, their token construction based on multi-channel values along the temporal dimension shares the same issues we highlight.
@@
Some recent methods (Liu et al., 2024a; Wang et al., 2024) transpose the input format above so each token represents an individual series, forming the Series-wise Transformer. This format, shown in Figure 2 (b), can be expressed as follows:
$\mathbf{Z}^{L\times C}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{C}]=\begin{bmatrix}\mathbf{P}_{1:C}\\ X_{I}^{\top}\\ \mathbf{O}\\ \end{bmatrix}=\begin{bmatrix}\mathbf{p}_{1}&X_{1}^{(1)}&\ldots&X_{1}^{(L_{I})}&o_{L_{I}+1}&\ldots&o_{L}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{p}_{C}&X_{C}^{(1)}&\ldots&X_{C}^{(L_{I})}&o_{L_{I}+1}&\ldots&o_{L}\\ \end{bmatrix}^{\top}$.
However, iTransformer (Liu et al., 2024a) introduced this formulation believing it enhances modeling of inter-series dependencies compared to the Temporal-wise Transformer, which contrasts with our ICL analysis. Each token in the Series-wise Transformer represents an entire input series, capturing relationships between timesteps within that series. For series $j$, the Transformer learns a predictor $f^{*}_{j}:X_{j}^{(1:L_{I})}\rightarrow X_{j}^{(L_{I}+1:L_{I}+L_{P})}$, which also adjusts its own parameters to the inputs $\{X^{(1:L_{I})}_{i}\}_{i\neq j}$ from other series. Essentially, it acts more like a univariate predictor capable of adapting its parameters based on the context series. It is capable of modeling temporally-aligned inter-series dependencies, but it is difficult to model shifted dependencies across timesteps.Similar to supervised learning, ICL treats each input within a token as independent features. This means that values across different timesteps are considered distinct features. The algorithm or model generated by ICL excels at building relationships within input features and can build relationship on the same feature (i.e. same timestep) across different data points (different series here) by performing linear combinations. However, it is less inclined to establish relationships between different features across different data points. In traditional supervised learning settings, different features from different data points are usually unrelated, so no relationship is built, but this contradicts the requirements of multivariate TSF settings when using this series-as-datapoint setting. We show in the last part of §3 that the Series-wise Transformer is better suited for real-world datasets with weak multivariate relationships than those with strong shifted inter-series dependencies, supporting our view from ICL. Also from this perspective, the Series-wise Transformer lacks ground truth context examples for the predictor, meaning it doesn’t fully utilize the Transformer’s ICL abilities. This limits its generalization and few-shot learning capabilities.
@@
Adaptive Model Reduction of ICTSP Existing complex Temporal-wise Transformers often struggle to outperform simple baselines, like linear predictors, in noisy datasets lacking clear temporal-wise and channel-wise dependencies. In contrast, ICTSP adapts to different dataset complexities and reduces to simpler models when necessary. Specifically: i) When temporal-wise dependencies are weak, ICTSP can adaptively select simpler functions for TSF, such as linear predictors and shallow MLPs, as achievable by ICL. ii) When series patterns change greatly over time where the impact of context examples is weak, ICTSP can reduce to a Series-wise Transformer focusing only on local lookback of target tokens without context tokens. iii) When local lookback from other series have minimal contribution, ICTSP degrades to a simple univariate MLP, in which case the inter-token information in attention term from Eq.equation 1 disappears, and token $j$ at layer $k$ becomes $\mathbf{z}_{j,k}=\mathbf{z}_{j,k-1}+\mathtt{LN}(\mathtt{FFN}_{j,k}(\mathbf{z}_{j,k-1}))$. If $\mathtt{FFN}$ also becomes ineffective, ICTSP processes target tokens only with input and output linear projections outside the Transformer. At this stage, the model reduces to a linear predictor, effective for highly noisy time series data. ICTSP’s design provides appropriate shortcuts to enhance adaptability. We further explain in §A.4.1 how inappropriate shortcuts in the model reduction of the Temporal-wise Transformer limit its ability to learn dynamic patterns.
@@ 2.3 Solving Key Issues of TSF Transformers
PatchTST (Nie et al., 2022) suggested that univariate Temporal-wise Transformers, which inputs each channel independently (Figure 3 (b)), can address overfitting. This approach breaks the series-wise relationship in the tokens, transforming the ICL predictor into $f^{*}_{j,t}:\mathbf{p}_{t}\rightarrow X^{(t)}_{j}$, which predicts values solely from positions. This shift enhances the model’s focus on temporal dependencies while ignoring series-wise dependencies, making the formulation more suitable for real-world TSF tasks. Notably, the prompt structure used by recent LLMs for TSF, "… timestep: value …", also benefits from this approach. Existing methods like ForecastPFN (Dooley et al., 2024) use similar structures.
ARM (Lu et al., 2024) introduced random dropping (Figure 3 (c)), which mitigates overfitting by randomly removing series from both input and output during training. This randomly breaks some relationships between channels within the tokens. With only positional information consistently preserved, the model focus more on the temporal generation process and retains only the most likely channel dependencies, reducing overfitting to non-existent channel relationships. However, the degree of channel dropout requires further tuning based on different dataset characteristics.
Methods such as LogTrans (Li et al., 2019) and ARM (Lu et al., 2024) use strategies similar to Figure 3 (d) to build local representations for tokens across time intervals instead of single timesteps. This introduces temporal relationships within the token. For TSF data with weak channel dependencies, tokens naturally focus on intra-series temporal relationships (solid lines in Figure 3 (d)) rather than the weak channel-wise relationships (dashed lines in Figure 3 (d)) as in the original token format. This approach alleviates overfitting but requires careful tuning of the local window selection. Recent methods like Liu et al. (2024b) tend to combine method (c) and (d) to further improve the generalization ability.
@@
Permutation Invariance Studies like DLinear (Zeng et al., 2023) and iTransformer (Liu et al., 2024a) suggest that the poor performance of Temporal-wise Transformers is partly due to the inherent permutation invariance issue in Transformer, meaning swapping the positions of input tokens does not significantly affect the output. The Temporal-wise Transformer treats context tokens more as representations of the underlying channel structure. Thus, when a dataset lacks a temporally varying underlying channel structure, $\mathbf{p}_{t}$ in $f^{*}_{t}:\mathbf{p}_{t}\rightarrow[X^{(t)}_{1},\cdots,X^{(t)}_{C}$$^{\top}$ becomes ineffective, and the output tends to converge to the context’s average. This is consistent with DLinear observations (Zeng et al., 2023) where shuffling of input token minimally impact the output. Conversely, in the ICTSP, each context token represents a forecasting training example, so swapping their positions does not harm the forecasting. This makes the ICTSP better suited to the Transformer’s characteristics for TSF compared to previous structures.
Channel Structure Restriction In the Temporal-wise Transformer, the number of input series and their relative positions within the tokens must be fixed. This limits the model’s ability to adapt to datasets with varying series numbers. Channel independence (Nie et al., 2022) may simplify the Temporal-wise Transformer to a univariate model, enhancing generalization but losing the ability to model inter-series dependencies. The Series-wise Transformer does not restrict the number of input series structurally but learns the characteristics of $C$ series during training through $\mathbf{p}_{j}$, making it hard to transfer embeddings to other datasets. Conversely, the ICTSP does not restrict the number of series and includes in-context training examples within the input, enabling easy transfer to input data of any size. Here, positional embeddings can be trained to only distinguish between samples without describing the characteristics of the series. The flexibility in input channel structure provides zero-shot multivariate forecasting ability, which previous multivariate models could not achieve.